# Optimal Online Change Detection via Random Fourier Features

### v1

@@ 1 Introduction
While algorithms for the problem of non-parametric online change point detection have been proposed—see Section 2 for a brief overview—state-of-the-art methods suitable for modern data suffer from at least one of the following two limitations. First, most procedures are not genuinely online from a statistical perspective, in the sense that they either assume the pre-change distribution to be known completely or they assume having access to historical data known to be from the pre-change distribution. Second, most approaches require the user to specify a window parameter over which local tests for a change in distribution will be applied. Choosing such a window is notoriously difficult [43], and choosing the window too small or too large leads to a reduction in power or an increase in the detection delay, respectively.
@@
  • Empirical validation: We perform a suite of benchmarks on synthetic data and on the MNIST data set to demonstrate the applicability of the proposed method. Our approach achieves competitive results throughout all experiments.
The article is structured as follows. We recall related work in Section 2 and introduce our notations and the problem in Section 3. Section 4 presents our algorithm and its guarantees, and Section 5 its minimax optimality. Experiments are in Section 6 and limitations in Section 7. Additional results and all proofs are in the supplement.
Table 1: Comparison of kernel-based change detectors. Genuinely online  — whether the algorithm can be executed without reference data known to be from the pre-change distribution; Window free  — whether the algorithm requires selection of a window parameter over which the detection statistic is calculated; Time comp. — runtime complexity per new observation; Space comp. — total space complexity; $n$ — total number of observations.
@@ 2 Related work
To tackle change point detection on multivariate data, many approaches exist; see Wang and Xie [54] for a survey. For example, Yilmaz [57], Kurt et al. [25] introduce procedures using summary statistics based on geometric entropy minimization, but assume knowledge of the pre-change distribution. An alternative non-parametric approach is using information on distances between data points [6, 7]. However, here one can construct alternatives which the procedures will always fail to detect as the limits of the test statistics employed do not metricize the space of probability distributions.
One principled approach to tackle this challenging setting is using kernel-based two sample tests via the MMD, which we recall briefly in Section 3.2. The key property, if the underlying kernel is characteristic [10, 51], is that the MMD metricizes the space of probability distributions and thus allows detecting any change. However, the sample estimators of the MMD typically have a quadratic runtime complexity, which prohibits their direct application in the online setting. To overcome this challenge, Zaremba et al. [61] propose Scan $B$-statistics, which achieve sub-quadratic time complexity by splitting the data into blocks and computing the quadratic-time MMD on each block. Consequently, Li et al. [29, 30] introduce an online change point detection algorithm that recursively estimates the Scan $B$-statistic over a sliding window. Extending this idea, Wei and Xie [55] propose an algorithm which performs the same operation over grid of windows of different sizes. However, these algorithms all require the choice of a window parameter and are not genuinely online as they require historical data known to be from the pre-change distribution. Keriven et al. [23] propose comparing two exponentially smoothed MMD statistics, where the MMD is approximated by using random Fourier features. However, the window selection problem is not avoided because, as shown by the authors, the smoothed statistic can be interpreted as computing differences between MMDs calculated on two windows of different sizes.
We summarize the kernel-based approaches coming with theoretical guarantees in Table 1 and note that we present an extension of our proposed Online RFF-MMD approach that allows taking historical data into account in the supplement. [^1] In Table 1, where applicable, $r$ denotes the number of random Fourier features, $W$ denotes the size of the window, and $N$ denotes the number of blocks of historical data. We treat the dimension of the data as fixed.
@@ 3 Preliminaries
In this section, we formally introduce the online change point detection problem (Section 3.1), and recall kernel-based two sample testing together with its RFF-based approximation (Section 3.2).
@@ 3.1 Problem statement
[…]
$H_{1,n}:\exists\eta<n\hskip 5.0pt\text{s.t}\hskip 5.0ptX_{t}\sim\begin{cases}\mathbb{P}&\text{ if }1\leq t\leq\eta\\ \mathbb{Q}&\text{ if }\eta<t\leq n\end{cases},\hskip 5.0pt\text{and some}\hskip 5.0pt\mathbb{P},\mathbb{Q}\in\mathcal{M}_{1}^{+}\hskip 5.0pt\text{where}\hskip 5.0pt\mathbb{P}\neq\mathbb{Q},$ (1)
@@ 3.2 Fast kernel-based two sample tests
To resolve the change point detection problem (3.1), we embed fast two sample tests based on RFF approximations to the MMD into a particular sequential testing scheme. In this section, we briefly recall the MMD statistic and its RFF approximation.
Let $\mathcal{H}_{K}$ be a reproducing kernel Hilbert space (RKHS; [2, 52]) on $\mathbb{R}^{d}$ with (reproducing) kernel $K:\mathbb{R}^{d}\times\mathbb{R}^{d}\to\mathbb{R}$. Denote by $\text{supp}(\Lambda)=\overline{\{A\in\sigma(\mathbb{R}^{d})\mid\Lambda(A)>0\}}$ the support of a measure $\Lambda$ on $\mathbb{R}^{d}$, where $\overline{A}$ denotes the closure of a set $A$. Throughout the article, we make the following assumption on the kernel:
@@
To simplify exposition, we also assume that $K(0,0)=1$, which can be achieved by scaling any bounded kernel. The conditions in Assumption 1 are satisfied by a number of commonly used kernels, including the Gaussian, Laplace, and B-spline kernels [51, Table 2].
@@
where the integral is meant in Bochner’s sense [9, Chapter II.2]. The continuity and boundedness assumptions ensure the existence of $\mu_{K}(\mathbb{P})$ for any $\mathbb{P}\in\mathcal{M}_{1}^{+}$ [51, Proposition 2]. Expression (4) gives rise to the MMD, which quantifies the distance between two measures $\mathbb{P},\mathbb{Q}\in\mathcal{M}_{1}^{+}$ via the distance between their mean embeddings in terms of the RKHS norm, and takes the form
@@
Crucially, the characteristic assumption ensures that (4) is injective and that the MMD metricizes the space $\mathcal{M}_{1}^{+}$ [51, Theorem 9], implying $\operatorname{MMD}_{K}[\mathbb{P},\mathbb{Q}]=0$ iff. $\mathbb{P}=\mathbb{Q}$. Given samples $X_{1},\ldots,X_{n}\stackrel{{\scriptstyle\text{i.i.d.}}}{{\sim}}\mathbb{P}$ and $Y_{1},\ldots,Y_{m}\stackrel{{\scriptstyle\text{i.i.d.}}}{{\sim}}\mathbb{Q}$ with associated empirical measures ^$\hat{\mathbb{P}}_{n}=:X_{1:n}$ and ^$\hat{\mathbb{Q}}_{m}=:Y_{1:m}$, respectively, the squared plug-in estimator of $\operatorname{MMD}_{K}[\mathbb{P},\mathbb{Q}]$ takes the form
$\left(\operatorname{MMD}_{K}\!\left[X_{1:n},Y_{1:m}\right]\right)^{2}=\left\|\mu_{K}\!\left(\hat{\mathbb{P}}_{n}\right)-\mu_{K}\!\left(\hat{\mathbb{Q}}_{n}\right)\right\|_{\mathcal{H}_{K}}^{2}=\left\|\frac{1}{n}\sum_{i=1}^{n}K\!\left(\cdot,X_{i}\right)-\frac{1}{m}\sum_{i=1}^{m}K\!\left(\cdot,Y_{i}\right)\right\|_{\mathcal{H}_{K}}^{2}$
@@
Random Fourier features [41, 50] alleviate this bottleneck in the offline setting; we recall the method in the following. For some $\omega\in\mathbb{R}^{d}$ write $\zeta_{\omega}(\mathbf{x})=e^{i\omega^{\mkern-1.5mu\mathsf{T}}\mathbf{x}}$, where $i=\sqrt{-1}$. By Bochner’s theorem,
$K\!\left(\mathbf{x},\mathbf{y}\right)=\psi(\mathbf{x}-\mathbf{y})=\int_{\mathbb{R}^{d}}e^{i\omega^{\mkern-1.5mu\mathsf{T}}(\mathbf{x}-\mathbf{y})}\mathrm{d}\Lambda\!\left(\omega\right)=\mathbb{E}_{\omega\sim\Lambda}\!\left[\zeta_{\omega}\left(\mathbf{x}\right)\zeta_{\omega}\left(\mathbf{y}\right)^{*}\right],$ (7)
[…]
@@ 4.1 The RFF-MMD stopping time
$N=\inf\left\{n\geq 2\mid\bigcup_{j=1}^{\left\lfloor\log_{2}(n)\right\rfloor-1}\sqrt{\frac{2^{j}(n-2^{j})}{n}}\operatorname{MMD}_{\hat{K}}\!\left[X_{1:(n-2^{j})},X_{(n-2^{j}+1):n}\right]>\lambda_{n}\right\},$ (10)
@@ 4.2 The RFF-MMD algorithm
We now present an efficient implementation of (10) and analyze its runtime and space complexity. We show the pseudo code of our proposed method in Algorithm 1; see also the corresponding figure in the supplement for a visual summary. The details are as follows. For each new observation $X_{t}$, we create a new window $W$, storing $z=\hat{z}_{K}(X_{t})$ and $c=1$ (Lines 56). The window $W$ is then added to the list of all windows $\mathcal{W}$ (Line 7). The remaining algorithm has two main parts.
@@
12:         if $\sqrt{\frac{c_{1}c_{2}}{c_{1}+c_{2}}}\operatorname{MMD}_{\hat{K}}\geq\lambda_{t}$ then
13:              print Change detected at element $\text{{print} Change detected at element }X_{t}\text{; most likely at position}i$.
14:              Drop tail of $\text{Drop tail of }\mathcal{W}$
@@ 4.3 Theoretical results
Next, we study the detection delay incurred by (10) when the threshold sequence is chosen to control the uniform false alarm probability at some level $\alpha\in(0,1)$. In the following we assume that the data take values in a compact subset of $\mathbb{R}^{d}$, and denote the Lebesgue measure of a set by ||$\cdot$|| ⋅ |. The following result shows that with high probability, provided the number of RFFs is chosen sufficiently large, the detection delay incurred by (10) is bounded from above by a quantity depending only on the chosen $\alpha$, the number of pre-change observations, and the squared MMD between the pre- and post-change distributions.
@@
$\left(N-\eta\right)^{+}\leq 1\vee C_{3}\frac{\log\left(2\eta/\alpha\right)}{\left(\operatorname{MMD}_{K}\!\left[\mathbb{P},\mathbb{Q}\right]\right)^{2}},$
where $C_{1}$, $C_{2}$, and $C_{3}$ are absolute constants independent of $\eta$, $\alpha$, and $\operatorname{MMD}_{K}\!\left[\mathbb{P},\mathbb{Q}\right]$, and, with $\sigma^{2}=\int\left\|\omega\right\|_{2}^{2}\mathrm{d}\Lambda\left(\omega\right)$, we have put
@@
Condition (11) can be interpreted as a signal strength requirement, measuring the strength according to the number of observations from the pre-change distribution and the squared MMD between $\mathbb{P}$ and $\mathbb{Q}$. The term $\log(2\eta/\alpha)$ reflects the cost of multiple testing when the data are drawn from $\mathbb{P}$. Such requirements are unavoidable from the minimax perspective in the corresponding offline problem [58, 38, 37], and the discussion in Yu et al. [59, Section 4.1] suggests that the same is true for genuinely online change point problems.
The requirement on the number of RFFs in (12) depends on $\operatorname{MMD}_{K}\!\left[\mathbb{P},\mathbb{Q}\right]$, which is unknown in practice. However, if one assumes an asymptotic setting with a fixed distance between $\mathbb{P}$ and $\mathbb{Q}$, and $\alpha\downarrow 0$, then (12) suggests that the number of RFFs should be chosen as $r=\Theta(\log 1/\alpha)$. To put this result into perspective, we compare it to online change procedures having a window. Here, the optimal window length also depends on the distance between the pre- and post-change distributions [30, 55]. However, choosing the window larger than this quantity can lead to an increase in the detection delay. This is not the case for RFF-MMD: in practice, one may choose the number of RFFs as large as possible subject to computational constrains, and choosing a larger number of RFFs does not negatively impact the detection delay.

### v2

@@ 1 Introduction
While algorithms for the problem of non-parametric online change point detection have been proposed—see Section 2 for a brief overview—state-of-the-art methods suitable for modern data suffer from at least one of the following two limitations. First, most procedures are not genuinely online from a statistical perspective, in the sense that they either assume the pre-change distribution to be known completely or they assume having access to historical data known to be from the pre-change distribution. Second, most approaches require the user to specify a window parameter over which local tests for a change in distribution will be applied. Choosing such a window is notoriously difficult [47], and choosing the window too small or too large leads to a reduction in power or an increase in the detection delay, respectively. Moreover, the window size imposes a limit on the historic data considered, rendering detecting sufficiently small changes impossible.
@@
  • Empirical validation: We perform a suite of benchmarks on synthetic data, the MNIST data, and the HASC data to demonstrate the applicability of the proposed method. Our approach achieves competitive results throughout all experiments.
The article is structured as follows. We recall related work in Section 2 and introduce our notations and the problem in Section 3. Section 4 presents our algorithm and its guarantees, and Section 5 its minimax optimality. Experiments are in Section 6 and limitations in Section 7. Additional results and all proofs are in the appendices.
Table 1: Comparison of kernel-based change detectors. Genuinely online — whether the algorithm can be executed without reference data known to be from the pre-change distribution; Window free — whether the algorithm requires selection of a window parameter over which the detection statistic is calculated; Time comp. — runtime complexity per new observation; Space comp. — total space complexity; $n$ — total number of observations. [^2] In Table 1, where applicable, $r$ denotes the number of random Fourier features, $W$ denotes the size of the window, and NN denotes the number of blocks of historical data. We treat the dimension of the data as fixed.
@@ 2 Related work
To tackle change point detection on multivariate data, many approaches exist; see Wang and Xie [58] for a survey. For example, Yilmaz [62], Kurt et al. [28] introduce procedures using summary statistics based on geometric entropy minimization, but assume knowledge of the pre-change distribution. An alternative non-parametric approach is using information on distances between data points [7, 8]. However, here one can construct alternatives which the procedures will always fail to detect as the limits of the test statistics employed do not metrize the space of probability distributions.
One principled approach to tackle this challenging setting is using kernel-based two sample tests via the MMD, which we recall briefly in Section 3.2. The key property, if the underlying kernel is characteristic [11, 55], is that the MMD metrizes the space of probability distributions and thus allows detecting any change. However, the sample estimators of the MMD typically have a quadratic runtime complexity, which prohibits their direct application in the online setting. To overcome this challenge, Zaremba et al. [66] propose Scan $B$-statistics, which achieve sub-quadratic time complexity by splitting the data into blocks and computing the quadratic-time MMD on each block. Consequently, Li et al. [32, 33] introduce an online change point detection algorithm that recursively estimates the Scan $B$-statistic over a sliding window. Extending this idea, Wei and Xie [59] propose an algorithm which performs the same operation over grid of windows of different sizes. However, these algorithms all require the choice of a window parameter and are not genuinely online as they require historical data known to be from the pre-change distribution. Keriven et al. [24] propose comparing two exponentially smoothed MMD statistics, where the MMD is approximated by using random Fourier features. However, the window selection problem is not avoided because, as shown by the authors, the smoothed statistic can be interpreted as computing differences between MMDs calculated on two windows of different sizes.
We summarize the kernel-based approaches coming with theoretical guarantees in Table 1 and note that we present two extensions of our proposed Online RFF-MMD approach in the appendix. The first allows taking historical data into account (Appendix A.1). The second permits detecting multiple change points within the same stream, with theoretical guarantees (Appendix A.2).
@@ 3 Preliminaries
In this section, we formally introduce the online change point detection problem (Section 3.1) and recall kernel-based two sample testing together with its RFF-based approximation (Section 3.2).
@@ 3.1 Problem statement
$H_{0,n}:X_{t}\sim\mathbb{P}\hskip 5.0pt\text{ for each }\hskip 5.0ptt\leq n\hskip 5.0pt\text{and some}\hskip 5.0pt\mathbb{P}\in\mathcal{M}_{1}^{+}\hskip 5.0pt\text{versus}$
$H_{1,n}:\exists\eta<n\hskip 5.0pt\text{s.t.}\hskip 5.0ptX_{t}\sim\begin{cases}\mathbb{P}&\text{ if }1\leq t\leq\eta\\ \mathbb{Q}&\text{ if }\eta<t\leq n\end{cases},\hskip 5.0pt\text{and some}\hskip 5.0pt\mathbb{P},\mathbb{Q}\in\mathcal{M}_{1}^{+}\hskip 5.0pt\text{where}\hskip 5.0pt\mathbb{P}\neq\mathbb{Q},$
@@ 3.2 Fast kernel-based two sample tests
To resolve the change point detection problem (Section 3.1), we embed fast two sample tests based on RFF approximations to the MMD into a particular sequential testing scheme. In this section, we briefly recall the MMD statistic and its RFF approximation.
Let $\mathcal{H}_{K}$ be a reproducing kernel Hilbert space (RKHS; [2, 56]) on $\mathbb{R}^{d}$ with (reproducing) kernel $K:\mathbb{R}^{d}\times\mathbb{R}^{d}\to\mathbb{R}$. Denote by $\text{supp}(\Lambda)=\overline{\{A\in\sigma(\mathbb{R}^{d})\mid\Lambda(A)>0\}}$ the support of a Borel measure $\Lambda$ on $\mathbb{R}^{d}$, where $\overline{A}$ denotes the closure of a set $A$ and $\sigma(\mathbb{R}^{d})$ the Borel $\sigma$-algebra on $\mathbb{R}^{d}$. Throughout the article, we make the following assumption on the kernel:
@@
To simplify exposition, we also assume that $K(\bm{0},\bm{0})=1$, which can be achieved by scaling any bounded kernel. The conditions in Assumption 1 are satisfied by several commonly used kernels, including the Gaussian kernel, mixtures of Gaussians, inverse multi-quadratic kernels, Matérn kernels, Laplace kernels, or B-spline kernels [55]. Assumption 1 permits to approximate the respective kernel function by finite-dimensional feature maps through Bochner’s theorem (elaborated below), which, in turn, allows the effective online estimation of MMD detailed in Section 4.
@@
where the integral is meant in Bochner’s sense [10, Chapter II.2]. The continuity and boundedness assumptions ensure the existence of $\mu_{K}(\mathbb{P})$ for any $\mathbb{P}\in\mathcal{M}_{1}^{+}$ [55, Proposition 2]. Expression (3) gives rise to the MMD, which quantifies the distance between two measures $\mathbb{P},\mathbb{Q}\in\mathcal{M}_{1}^{+}$ via the distance between their mean embeddings in terms of the RKHS norm, and takes the form
@@
Crucially, the characteristic assumption ensures that (3) is injective and that the MMD metrizes the space $\mathcal{M}_{1}^{+}$ [55, Theorem 9], implying $\operatorname{MMD}_{K}[\mathbb{P},\mathbb{Q}]=0$ iff. $\mathbb{P}=\mathbb{Q}$. Given samples $X_{1},\ldots,X_{n}\stackrel{{\scriptstyle\text{i.i.d.}}}{{\sim}}\mathbb{P}$ and $Y_{1},\ldots,Y_{m}\stackrel{{\scriptstyle\text{i.i.d.}}}{{\sim}}\mathbb{Q}$ with associated empirical measures $\hat{\mathbb{P}}_{n}=:X_{1:n}$ and $\hat{\mathbb{Q}}_{m}=:Y_{1:m}$, respectively, the squared plug-in estimator of $\operatorname{MMD}_{K}[\mathbb{P},\mathbb{Q}]$ takes the form
$\left(\operatorname{MMD}_{K}\!\left[X_{1:n},Y_{1:m}\right]\right)^{2}=\left\|\mu_{K}\!\left(\hat{\mathbb{P}}_{n}\right)-\mu_{K}\!\left(\hat{\mathbb{Q}}_{n}\right)\right\|_{\mathcal{H}_{K}}^{2}\hskip-2.84544pt=\left\|\frac{1}{n}\sum_{i=1}^{n}K\!\left(\cdot,X_{i}\right)-\frac{1}{m}\sum_{i=1}^{m}K\!\left(\cdot,Y_{i}\right)\right\|_{\mathcal{H}_{K}}^{2}$
@@
Random Fourier features [45, 54] alleviate this bottleneck in the offline setting; we recall the method in the following. For some $\omega\in\mathbb{R}^{d}$ write $\zeta_{\omega}(\mathbf{x})=e^{-i\omega^{\mkern-1.5mu\mathsf{T}}\mathbf{x}}$, where $i=\sqrt{-1}$. By Bochner’s theorem,
$K\!\left(\mathbf{x},\mathbf{y}\right)=\psi(\mathbf{x}-\mathbf{y})=\int_{\mathbb{R}^{d}}e^{-i\omega^{\mkern-1.5mu\mathsf{T}}(\mathbf{x}-\mathbf{y})}\mathrm{d}\Lambda\!\left(\omega\right)=\mathbb{E}_{\omega\sim\Lambda}\!\left[\zeta_{\omega}\left(\mathbf{x}\right)\zeta_{\omega}\left(\mathbf{y}\right)^{*}\right],$
where denotes the complex conjugate and, using that $K(0,0)=1$, one has that $\Lambda\in\mathcal{M}_{1}^{+}$. As $\Lambda$ and $K$ are real-valued, Euler’s identity implies that $\int e^{-i\omega^{\mkern-1.5mu\mathsf{T}}(\mathbf{x}-\mathbf{y})}\mathrm{d}\Lambda\!\left(\omega\right)=\int\cos\left(\omega^{\mkern-1.5mu\mathsf{T}}\!\left(\mathbf{x}-\mathbf{y}\right)\right)\mathrm{d}\Lambda\!\left(\omega\right)$. Therefore, using that $\cos(\alpha-\beta)=\cos\alpha\cos\beta+\sin\alpha\sin\beta$, picking some $r\in\mathbb{N}$, and sampling $\omega_{1},\dots,\omega_{r}\overset{\text{i.i.d.}}{\sim}\Lambda$, a low variance estimator for $K(\mathbf{x},\mathbf{y})$ is given by
@@ 4.1 The RFF-MMD stopping time
$N=\inf\left\{n\geq 2\mid\bigcup_{j=0}^{\left\lfloor\log_{2}(n)\right\rfloor-1}\sqrt{\frac{2^{j}(n-2^{j})}{n}}\operatorname{MMD}_{\hat{K}}\!\left[X_{1:(n-2^{j})},X_{(n-2^{j}+1):n}\right]>\lambda_{n}\right\},$ (8)
@@ 4.2 The RFF-MMD algorithm
We now present an efficient implementation of (8) and analyze its runtime and space complexity. We show the pseudo code of our proposed method in Algorithm 1; see also Example 1 and Figure 1 for a summary. The details are as follows. For each new observation $X_{t}$, we create a new window $W$, storing $z=\hat{z}_{K}(X_{t})$ and $c=1$ (Lines 56). The window $W$ is then added to the list of all windows $\mathcal{W}$ (Line 7). The remaining algorithm has two main parts.
@@
12:   if $\sqrt{\frac{c_{1}c_{2}}{c_{1}+c_{2}}}\operatorname{MMD}_{\hat{K}}>\lambda_{t}$ then
[…]
14:     return
@@
The following Example 1 and the corresponding Figure 1 illustrate how Algorithm 1 operates upon observing the first $6$ samples $X_{1},\ldots,X_{6}$.
Example 1.
When observing the first element $X_{1}$, the algorithm creates a new window $W$, storing the feature map $\hat{z}_{K}(X_{1})$ and that $W$ has one element ($c=1$). Similarly, when observing $X_{2}$, the algorithm creates a new window $W^{\prime}$, storing $\hat{z}_{K}(X_{2})$ and $c=1$. As both windows, $W$ and $W^{\prime}$, have the same counts, the algorithm merges them into a new window $W$, storing $\hat{z}_{K}(X_{1})+\hat{z}_{K}(X_{2})$ and $c=2$, thereby maintaining the dyadic structure. The algorithm proceeds in this manner, resulting in the construction of the dyadic grid outlined in Section 4.1. For example, when observing $X_{6}$, there are again two windows of size $1$, which the algorithm merges to store a total of two windows, one capturing $X_{1},\ldots,X_{4}$ and the other one capturing $X_{5},X_{6}$. We recall that the observations themselves are never stored explicitly.
Before presenting out theoretical results in the following section, we emphasize that Algorithm 1 matches (8), that is, there is no approximation error.
@@ 4.3 Theoretical results
Next, we study the detection delay incurred by (8) when the threshold sequence is chosen to control the uniform false alarm probability at some level $\alpha\in(0,1)$. In the following, we assume that the data take values in a compact subset of $\mathbb{R}^{d}$, and denote the Lebesgue measure of a set by $|\cdot|$. The following result shows that with high probability, provided the number of RFFs is chosen sufficiently large, the detection delay incurred by (8) is bounded from above by a quantity depending only on the chosen $\alpha$, the number of pre-change observations, and the squared MMD between the pre- and post-change distributions.
@@
$\left(N-\eta\right)^{+}\leq 1\vee C_{3}\frac{\log\left(2\eta/\alpha\right)}{\left(\operatorname{MMD}_{K}\!\left[\mathbb{P},\mathbb{Q}\right]\right)^{2}},$ (11)
where $C_{1}$, $C_{2}$, and $C_{3}$ are absolute constants independent of $\eta$, $\alpha$, and $\operatorname{MMD}_{K}\!\left[\mathbb{P},\mathbb{Q}\right]$, and, with $\sigma^{2}=\int_{\mathbb{R}^{d}}\left\|\omega\right\|_{2}^{2}\mathrm{d}\Lambda\left(\omega\right)$, we have put
@@
Condition (9) can be interpreted as a signal strength requirement, measuring the strength according to the number of observations from the pre-change distribution and the squared MMD between $\mathbb{P}$ and $\mathbb{Q}$. The term $\log(2\eta/\alpha)$ reflects the cost of multiple testing when the data are drawn from $\mathbb{P}$. Such requirements are unavoidable from the minimax perspective in the corresponding offline problem [63, 42, 41], and the discussion in Yu et al. [64, Section 4.1] suggests that the same is true for genuinely online change point problems.
The requirement on the number of RFFs in (10) depends on $\operatorname{MMD}_{K}\!\left[\mathbb{P},\mathbb{Q}\right]$, which is unknown in practice. However, if one assumes an asymptotic setting with a fixed distance between $\mathbb{P}$ and $\mathbb{Q}$, and $\alpha\downarrow 0$, then (10) suggests that the number of RFFs should be chosen as $r=\Theta(\log 1/\alpha)$. To put this result into perspective, we compare it to online change procedures having a window. Here, the optimal window length also depends on the distance between the pre- and post-change distributions [33, 59]. However, choosing the window larger than this quantity can lead to an increase in the detection delay. This is not the case for RFF-MMD: in practice, one may choose the number of RFFs as large as possible subject to computational constraints, and choosing a larger number of RFFs does not negatively impact the detection delay.